Louis and William were both working hard to lose weight. Louis weighed 330 pounds and William weighed 280 pounds. After 6 month, Louis had lost 40 pounds and William had lost 30 pounds. Who was having a bigger impact on his weight over this time period?
Answer by Theo(13342)

louis started at 330 and lost 40.
william started at 280 and lost 30.

louis wound up being (330 - 40) / 330 = .8787878788 of his original weight.

william wound up being (280 - 30) / 280 = .8928571429 of his original weight.

louis lost a bigger percentage of his total weight than william did.

40/330 = .1212121212 = 12.12% = the percent weight loss for louis.

30/280 = .1071428571 = 10.71% = the percent weight loss for william.